12. Solve the recurrence relation an=an-1+n2,where a0=7 by substitution


1
Expert's answer
2022-03-07T04:12:05-0500

Let us solve the recurrence relation "a_n=a_{n-1}+n^2," where "a_0=7."


The characteristic equation "k-1=0" has the root "k=1," and hence the general solutionis of the form:

"a_n=C\cdot 1^n+b_n," where "b_n=n(an^2+bn+c)=an^3+bn^2+cn."

It follows that

"an^3+bn^2+cn=a(n-1)^3+b(n-1)^2+c(n-1)+n^2."

For "n=0" we get "0=-a+b-c," and hence "a=b-c."

For "n=1" we get "a+b+c=1," and thus "b-c+b+c=1," that is "b=\frac{1}2." Therefore, "a=\frac{1}2-c."

For "n=-1" we get "-a+b-c=-8a+4b-2c+1."

It follows that

"c-\frac{1}2+\frac{1}2-c=8c-4+2-2c+1," and therefore "6c=1."

We conclude that "c=\frac{1}6,\ a=\frac{1}2-\frac{1}6=\frac{1}3."

Consequently, the general solutionis of the form:

"a_n=C+\frac{1}3n^3+\frac{1}2n^2+\frac{1}6n."

It follows that "7=a_0=C," and we conclude that the particular solution is the following:

"a_n=7+\frac{1}3n^3+\frac{1}2n^2+\frac{1}6n."